# Generalized Multi-Objective Reinforcement Learning with Envelope Updates in URLLC-enabled Vehicular Networks

### v2

@@ V-B Baselines and Evaluation Metrics
To comprehensively evaluate the performance of MO-DDQN-envelope, we implement the followsing baseline algorithms for comparison: MO-DQN, MO-DDQN (as discussed in Section IV-A), MO-dueling-DDQN, and MO-PPO. MO-dueling-DDQN and MO-PPO are variations of MO-DDQN, utilizing single-policy multi-objective dueling DDQN and Proximal Policy Optimization (PPO) algorithms, respectively, instead of DDQN. These algorithms, specifically developed for performance evaluation, are extensions of single-policy RL methods originally proposed in [42] and [43], respectively. Additionally, we incorporate several state-of-the-art algorithms from MORL-Baselines [40] for comparison. These include Scalarized Q-learning for single-policy MORL (MO-Q) [44], Scalarized DQN for single-policy DQN (MO-DQN) [1], Action Branching Architectures with Dueling DDQN for MORL (MO-Dueling-DDQN) [11], Multi-objective exploration for proximal policy optimization (MO-PPO) [45], and Expected Utility Policy Gradient (EUPG) [46]. To evaluate the performance, we consider the following metrics. Assume episode $e$ ends on time step $T_{e}$. we define (1) total transportation reward: $R^{tran}_{e}=\mathbb{E}_{j\in M_{1}}[\sum_{t=1}^{T_{e}}r^{\mathrm{j,tran}}_{t}]$. (2) total telecommunication reward: $R^{tele}_{e}=\mathbb{E}_{j\in M_{1}}[\sum_{t=1}^{T_{e}}r^{\mathrm{j,tele}}_{t}]$. (3) collision rate: $\delta_{e}=1-\frac{T_{e}}{T_{hl}}$ (4) HOs Probability: $\xi_{e}=\mathbb{E}_{j\in M_{1}}[\xi^{j}_{T_{e}}]$. where $r^{\mathrm{j,tran}}_{t}$ and $r^{\mathrm{j,tele}}_{t}$ are instantaneous transportation and telecommunication rewards specified in equations (17) and (18), respectively. $T_{hl}$ represents the horizon limit of each episode. $\xi^{j}_{T_{e}}$ denotes HOs probability by the end of each episode ($T_{e}$’s step) defined in equation (18). Table IV evaluate total transportation reward (17), total telecommunication reward (18), total rewards defined by $R^{total}_{e}=R^{tele}_{e}+R^{tran}_{e}$ over MO-DQN, MO-DDQN, and MO-DDQN-Envelope.
Desire Velocity Model Number of AVs $M$
@@ V-C6 Training and Evaluation Time Complexity Per Step
  • Telecommunication Only This scenario considers only the telecommunication reward in the MDP design. The telecommunication reward has no impact on the training process.
  • Transportation Only This scenario focuses solely on the transportation reward, with the telecommunication reward not influencing the training.
  • Telecommunication+Transportation This scenario incorporates both the transportation and Telecommunication rewards jointly during the training process.
@@ V-C7 Simultaneously Multi Objective Optimization Analysis
As depicted in Figure 8, we consider the following three scenarios to examine the impact of optimizing different objectives. we found that considering both telecommunication and transportation objectives together performs well compared to relying on Telecommunication Only and Transportation Only alone.
@@ VI Conclusion
We introduce a novel MORL framework tailored for devising joint network selection and autonomous driving policies within a multi-band VNet. Our goals encompass enhancing traffic flow, minimizing collisions, maximizing data rates, and minimizing handoffs (HOs). We achieve this through controlling vehicle motion dynamics and network selection, employing a unique reward function that optimizes data rate, traffic flow, load balancing, and penalizes HOs and unsafe driving behaviors. The problem is formalized as a MOMDP, integrating telecommunication and autonomous driving utilities in its rewards. We propose single policy MORL solutions with predefined preferences, transforming the MOOP into a single-objective one, utilizing DQN and double DQN solutions to derive optimal policies dependent on relative preferences.
Addressing the challenge of learning optimized policies across varied preferences, we develop an envelope MORL solution. This approach enables effective navigation across preference spaces, generating tailored policies. Our algorithm leverages the contraction properties of the optimality operator governing a generalized Bellman equation and optimizes the convex envelope of multi-objective Q-values. Hindsight experience replay and homotopy optimization aid in manageable learning across diverse preferences. Additionally, we construct a novel simulation testbed, ”RF-THz-Highway-Env,” based on ”highway-env,” emulating a multi-band wireless network-enabled VNet. Numerical results demonstrate the superiority of our proposed solution over weighted sum-based MORL solutions with DQN, showcasing improvements of 12.7%, 18.9%, and 12.3% on average transportation reward, average telecommunication reward, and average HO rate, respectively.
Future research could extend the generalization of the MORL model, which is necessary to adapt to the dynamics of the traffic flow.

### v3

@@ V-B Baselines and Evaluation Metrics
To comprehensively evaluate the performance of MO-DDQN-envelope, we implement the followsing baseline algorithms for comparison: MO-DQN, MO-DDQN (as discussed in Appendix), MO-dueling-DDQN, and MO-PPO. MO-dueling-DDQN and MO-PPO are variations of MO-DDQN, utilizing single-policy multi-objective dueling DDQN and Proximal Policy Optimization (PPO) algorithms, respectively, instead of DDQN. These algorithms, specifically developed for performance evaluation, are extensions of single-policy RL methods originally proposed in [43] and [44], respectively. Additionally, we incorporate several state-of-the-art algorithms from MORL-Baselines [41] for comparison. These include Scalarized Q-learning for single-policy MORL (MO-Q) [45], Scalarized DQN for single-policy DQN (MO-DQN) [1], Action Branching Architectures with Dueling DDQN for MORL (MO-Dueling-DDQN) [12], Multi-objective exploration for proximal policy optimization (MO-PPO) [46], and Expected Utility Policy Gradient (EUPG) [47]. To evaluate the performance, we consider the following metrics. Assume episode $e$ ends on time step $T_{e}$. we define (1) total transportation reward: $R^{tran}_{e}=\mathbb{E}_{j\in M_{1}}[\sum_{t=1}^{T_{e}}r^{\mathrm{j,tran}}_{t}]$. (2) total telecommunication reward: $R^{tele}_{e}=\mathbb{E}_{j\in M_{1}}[\sum_{t=1}^{T_{e}}r^{\mathrm{j,tele}}_{t}]$. (3) collision rate: $\delta_{e}=1-\frac{T_{e}}{T_{hl}}$ (4) HOs Probability: $\xi_{e}=\mathbb{E}_{j\in M_{1}}[\xi^{j}_{T_{e}}]$. where $r^{\mathrm{j,tran}}_{t}$ and $r^{\mathrm{j,tele}}_{t}$ are instantaneous transportation and telecommunication rewards specified in equations (16) and (17), respectively. $T_{hl}$ represents the horizon limit of each episode. $\xi^{j}_{T_{e}}$ denotes HOs probability by the end of each episode ($T_{e}$’s step) defined in equation (17). Table IV evaluate total transportation reward (16), total telecommunication reward (17), total rewards defined by $R^{total}_{e}=R^{tele}_{e}+R^{tran}_{e}$ over MO-DQN, MO-DDQN, and MO-DDQN-Envelope.
Target Velocity Model Number of AVs $M$
@@ V-C6 Training and Evaluation Time Complexity Per Step
  • Telecommunication Only: This scenario considers only the telecommunication reward in the MDP design. The telecommunication reward has no impact on the training process.
  • Transportation Only: This scenario focuses solely on the transportation reward, with the telecommunication reward not influencing the training.
  • Telecommunication+Transportation: This scenario incorporates both the transportation and telecommunication rewards jointly during the training process.
@@ V-C7 Simultaneously Multi Objective Optimization Analysis
As depicted in Figure 8, we consider the three scenarios to examine the impact of optimizing different objectives. we found that considering both telecommunication and transportation objectives performs well compared to relying on Telecommunication Only and Transportation Only alone.
@@ VI Conclusion
We introduce a novel MORL framework tailored for devising joint network selection and autonomous driving policies within a multi-band VNet. Our goals encompass enhancing traffic flow, minimizing collisions, maximizing data rates, and minimizing handoffs (HOs). We achieve this through controlling vehicle motion dynamics and network selection, employing a unique reward function that optimizes data rate, traffic flow, load balancing, and penalizes HOs and unsafe driving behaviors. The problem is formalized as a MOMDP, integrating telecommunication and autonomous driving utilities in its rewards. We propose single policy and multi-policy MORL solutions with predefined and unknown preferences. Numerical results demonstrate the superiority of our proposed solution over weighted sum-based MORL solutions with DQN, showcasing improvements of 12.7%, 18.9%, and 12.3% on average transportation reward, average telecommunication reward, and average HO rate, respectively. Future research could enhance the generalization of the MORL model to better adapt to dynamic traffic conditions using well-established strategies such as meta-learning [50].
Acknowledgment
The authors would like to thank Dr. Hongda Wu for his valuable insights during the initial discussions and his contributions to this work’s development and improvement.
Appendix A Appendix
Given a set of preferences in MORL problems, single policy algorithms aim to scalarize the reward value to determine the best policy, considering the relative priorities assigned to competing objectives. We explore two DRL methods: DQN and DDQN for MORL, each of which employs a neural network parameterized by $\boldsymbol{\theta}_{t}$ (in each time step $t$) to approximate the $Q$-value function for a state-action pair, i.e., $Q(s_{t},a_{t};\boldsymbol{\theta}_{t})$. After taking action $a_{t}$ in state $s_{t}$ and receiving instant reward $r_{t+1}$, we can formulate a target $Q$ function,
${\hat{Q}}(s_{t},a_{t})=\mathbb{E}_{\pi}\left[r_{t}(s_{t},a_{t})+\mathcal{\gamma}Q_{\pi}(s_{t-1},a_{t-1})\right]$ (26)
which is used to optimize the neural network $\boldsymbol{\theta}_{t}$ using gradient descent, as given in [51],
The MO-DQN method incorporates a target Q-network and experience replay to stabilize the learning process and ensure convergence, as discussed in the following:
  • Experience Replay: To address issues related to correlations between sequential observations and to improve data efficiency, MO-DQN utilizes the experience replay mechanism, which stores past transition tuples $(s_{z},a_{z},s_{z+1},r_{z})$ in a replay buffer $\mathcal{D}_{\mathcal{T}}$ with size $N_{\mathcal{T}}$, i.e., $z\in\{1,\dots,N_{\mathcal{T}}\}$. During the training phase, mini-batches of these transitions are randomly sampled from the buffer. This method not only reduces the variance of each update but also allows the neural network to benefit from learning across a diverse range of past experiences, thus avoiding local optima and overfitting.
Initialization:
Initialize $Q$-network weights $\boldsymbol{\theta}$ randomly,
while episode $<$ episode limit and runtime $<$ time limit do
       Initialize $t\leftarrow 0$ and state $s_{t}$ based on environment
       while $t\leq T_{hl}$  do
             Derive $a^{\mathrm{tran}}_{t}$ and $a^{\mathrm{tele}}_{t}$ from $a_{t}$
             Apply $a^{\text{tran}}_{t}$ and $a^{\text{tele}}_{t}$,
             observe reward $r_{t}$ and next state $s_{t+1}$.
             Store transition $(s_{t},a_{t},s_{t+1},r_{t})$ in $\mathcal{D}_{\mathcal{T}}$.
             where $z\in\{1,\ldots,N_{\mathcal{T}}\}$.
             Set target-$Q$ for each sampled transition:
             for each transition $z$ do
                   if episode ends at step $z+1$ then
                         $\hat{Q}(s_{z},a_{z})=r_{z}$
                   else
                   end if
             end for
             if MO-DDQN then
             end if
            $t\leftarrow t+1$
       end while
      Update policy $\pi$ based on learned $Q$.
end while
Algorithm 2 MO-DDQN Algorithm